If x+1/y=1 and y+1/z=1 then find the value of z+1/x

Grade:8

1 Answers

Arun
6 years ago
Dear student
Let assume for x = ½
Now
Y will be 2 from 1st equation.
Hence from 2nd equation,
1/z = 1-2 = -1
Hence z = -1
Now 
z + 1/x = -1 + 2 = 1
 
Regards
Arun(askIITians forum expert)